### v1

@@ 8.4. Fast Termination
Theorem 8.6 ().
FICC satisfies fast termination.
@@
With the additional assumption that $n-p$ replicas behave honestly during a phase of synchrony, we can guarantee $n-p$ fast shares for the rank-0 block reaching every replica. Thus, after a single round-trip time, each replica will be able to FP-finalize the rank-0 block. ∎
@@ 9.1. Methodology
As mentioned in section 8.1, a scenario with a different block proposal timing between FICC and ICC can only be crafted by reordering some of the messages exchanged between replicas. Therefore, in all experimental scenarios that address normal network conditions or enforce message order (e.g. use of TCP), FICC and ICC will exhibit identical behavior and timing with respect to dissemination and notarization of blocks. While FICC will of course behave differently to finalize received blocks, throughput is determined entirely by block creation interval, which remains unchanged.
We conclude that the throughput performance is unchanged in FICC  and bigger block payloads have the same effects on throughput and latency in FICC and ICC. To isolate the effects of the changes between the two protocols, we focus the experiments on the finalization latency of small blocks.
Application subnets of the Internet Computer feature 13 replicas (Dfinity, 2023a). To test the potential of FICC as a drop-in replacement for ICC, and since $n=16$ is optimal for both $f=5$, $p=0$ and $f=3$, $p=3$ experimental scenarios, we perform experiments with up to 16 replicas around the globe (see fig. 6). Replicas run on AWS t3.medium EC2 instances (with 2vCPUs, 4 GB of memory, running Ubuntu 22.04).
@@
Figure 6. The coordinates of the 16 deployed EC2 instances forming a global network are marked by black circles (section 9.3). The red triangles mark 13 clustered replicas and 3 eccentric replicas (section 9.4), while the blue squares show the coordinates of 4 distant replicas (section 9.5).
We wish to demonstrate that (i) in any network topology FICC is performing consistently as well as or better than ICC, while (ii) outperforming ICC in evenly geo-distributed settings. Finally, we want to show (iii) the positive effect of setting $p>0$, allowing for latency improvements even in the face of bottlenecks such as eccentric nodes.
As the goal of our evaluation is to measure the precise impact of our changes, we define latency as the proposal finalization time, both at the proposer (rank-0 replica) and at the other replicas. While the proposer finalization time can be measured precisely using the proposer’s system clock, other replicas have no information about when a round was started. We circumvent this inherent limitation, by relying on clock accuracy across AWS datacenters, which is handled by periodic synchronization to an NTP server (Amazon, 2023). This allows us to use the block timestamp as a round’s starting time.
@@
As required by the protocol, we set the proposal ($\Delta_{prop}$) and notarization ($\Delta_{notary}$) delays to be larger than the message delay experienced without network disruptions, such that our experimental results correspond to regular network conditions and only one block is proposed per round. Each experiment is run for 100 seconds, which is shown to be sufficient, as the measurements show remarkable periodicity (e.g., see fig. 7(b)).
9.2. Implementation
We implement FICC in Rust by adapting the existing implementation of the Internet Computer Protocol (Dfinity, 2023b) and open-source our code (Massimo Albarello, 2023). To test performance we strip the original codebase of all but its consensus layer and use libp2p’s flood-sub implementation (Protocol Labs, 2023) as the network layer. For clarity and simplicity, we also replace random beacon leader election with round-robin rotation.
@@ 9.3. Global Network
(a) Proposal finalization times across all replicas, for ICC and two FICC variants.
(b) FICC with f=5, p=0. Proposal finalization time at replica in Tokyo (ap-notheast-1). Own proposals are marked in red.
(c) FICC with f=5, p=0. Proposal finalization time at respective proposer.
Figure 7. Evaluation results of a network consisting of 16 globally distributed replicas.
We simulate a worldwide deployment akin to an IC subnet by distributing 16 replicas as shown by the black circles in fig. 6.
In this setting the average proposal finalization time measured for the ICC implementation was 845ms. By running FICC with $f=5,p=0$, an average proposal finalization time of 705ms is measured instead, while FICC with $f=3,p=3$ further reduces it to 638ms (fig. 6(a)).
Setting $f=5$ and $p=0$ leads to better results compared to ICC, as most replicas can FP-finalize blocks. In fig. 6(b), the proposal finalization time is shown from the point of view of the replica situated in Tokyo. Interestingly, one can observe that the fast path is almost always taken, except for its own proposals (marked by red dotted lines). Intuitively, this can be explained by the fact that in a highly connected network topology such as this one, fast shares from replicas situated between the proposer and recipient arrive just after the block proposal. (We hypothesize that this would not happen in e.g. a star network topology.)
Figure 6(c) shows that for some rounds IC-finalization (i.e., receiving two quorums of $n-f$ shares each) was still faster than waiting for a single quorum of $n-p$ fast shares. This is to be expected, especially for proposers on the periphery. Note also that fig. 6(a) shows a smaller variance in latency for ICC. This can be explained as ICC only waits for the $n-f=11$ fastest responses, while FICC requires $n-p$ shares to utilize the fast path (16 and 13 shares respectively). Thus, a remote node is expected to benefit less from the fast path. On the other hand, a node close to the center of the network will benefit more from the reduced number of communication steps as shown in fig. 6(a).
(a) Proposal finalization times across all replicas, for ICC and two FICC variants.
(b) FICC with f=5, p=0. Proposal finalization time as measured by the respective proposer.
(c) FICC with f=3, p=3. Proposal finalization time as measured by the respective proposer.
(d) FICC with f=5, p=0. Proposal finalization time across replicas.
(e) FICC with f=3, p=3. Proposal finalization time across replicas.
Figure 8. Evaluation results of a network consisting of 16 replicas, with 3 eccentric replicas.
9.4. Global Network with Eccentric Replicas
In this experiment, we demonstrate the effect of outliers (i.e., eccentric nodes) on the performance of FICC. As in section 9.3, 16 nodes are globally distributed, however, all except for three nodes are clustered in North America and Europe (see the red triangles in fig. 6). As fig. 7(a) shows, ICC averages a proposal finalization time of 853ms, which FICC with $p=0$ only marginally improves on (avg. finalization time of 844ms). This is to be expected, as the clustered nodes would need to wait for all three outliers to finalize a block through the fast path. Instead, they can IC-finalize blocks amongst themselves, without input needed from the outliers.
When the parameter $p$ is raised to $p=3$, the closely connected replicas can finalize blocks through the fast path by themselves, which explains the 20.4% drop in latency they experience in this scenario (679ms finalization time). Figure 7(d) and fig. 7(e) confirm that three replicas experience high latency. Even when $p=3$, their latency isn’t improved, however, the large cluster vastly benefits, as they now can finalize through the fast path, without having to wait for finalization shares from the eccentric replicas.
9.5. Global Network with Few Distant Replicas
Figure 9. Proposal finalization times across all replicas, for ICC and FICC respectively, obtained by benchmarking a network of 4 replicas for 100 seconds.
In this experiment, we measure the potential for FICC improvements under the adversity of an almost maximal message delay in a small network. Four replicas are distributed as shown by the blue squares in fig. 6. Due to the network’s size $p=0$ is forced. As fig. 9 shows, ICC averages a proposal finalization time of 1.30 seconds, which FICC manages to improve by 15.1% to 1.10 seconds. In absolute terms, this is a much higher latency than the previous experiments and shows that geolocation plays a major role when comparing algorithms in practice.
@@ 10. Conclusion
The Internet Computer Consensus protocol family borrows techniques from blockchains in order to improve the complexity of classical BFT protocols. The proposed Fast Internet Computer Consensus improves upon the original algorithm by providing a simultaneous dual mode, without incurring extra cost, thus closing the gap between state-of-the-art SMR protocols and classical fast consensus literature.
Concretely, we show that it is possible to achieve optimally fast proposal finalization time for a rotating leader SMR protocol. We experimentally demonstrate the effectiveness of the fast path by showing latency improvements compared to ICC. Ultimately, the effectiveness of the fast path depends largely on the network topology, as well as the parameter $p\in[0,f]$. If $p=0$, then a single corrupt replica can prevent fast path finalization by not sending its fast share. In the context of permissioned blockchains though, this setting is compelling, since it is possible to detect adversarial behavior and remove such misbehaving replicas (Buterin and Griffith, 2017; Camenisch et al., 2022). Moreover, in case a fast path robust to adversarial behavior or straggling replicas is desired, $p$ can be set arbitrarily close to $f$. This allows to take advantage of FICC even in the setting of persistent adversity and unfavorable network conditions.

### v2

@@ 8.4. Fast Termination
Theorem 8.7.
Banyan satisfies fast termination.
@@
With the additional assumption that $n-p$ replicas behave honestly during a phase of synchrony, we can guarantee $n-p$ fast votes for the rank-0 block reaching every replica. Thus, after a single round-trip time, each replica will be able to FP-finalize the rank-0 block. ∎
@@ 9. Evaluation
9.1. Implementation
We implement Banyan in Golang, utilizing the Bamboo BFT framework, which is designed for prototyping and evaluating chained BFT algorithms (Gai et al., 2021). This framework already supports the two well-known protocols, Streamlet (Chan and Shi, 2020) and HotStuff (Yin et al., 2019).
To enhance the performance and reliability of the framework, we introduced a few modifications. For instance, by using non-blocking message-passing channels internally, and by forwarding blocks that extend the tip of the chain, we drastically improve the performance of all algorithms implemented with Bamboo. The impact of the latter change is especially surprising to us, as originally only Streamlet pursued this strategy in the provided implementation, leading to an unexpected advantage over all other protocols, both in terms of throughput and latency. Henceforth, we have been careful to treat all protocols equally.
To further increase predictability and transparency of the protocol operations, we have replaced the random beacon leader election with a round-robin rotation. Our code and benchmarking scripts are made openly available (Sliwinski and Vonlanthen, 2024).
@@ 9.2. Methodology
Figure 5. The red triangles mark the coordinates of the 4 AWS datacenters used in Section 9.3, while the yellow crosses show the coordinates of 4 datacenters in the US used in Section 9.4. The 19 different coordinates of the deployed EC2 instances from Section 9.5 are marked by black circles.
Application subnets of the Internet Computer feature 13 replicas (Dfinity, 2024a). To test the potential of Banyan as a drop-in replacement for ICC, and since $n=19$ is optimal for both $f=6$, $p=1$ and $f=4$, $p=4$ experimental scenarios, we perform experiments with up to 19 replicas around the globe (see Figure 5). Replicas run on a testbed comprised of AWS t3.large EC2 instances (with 2vCPUs, 8 GB of memory, running Ubuntu 22.04).
We wish to demonstrate that (i) requiring only two rounds of communication instead of three has a positive impact and does not increase variance, while (ii) not suffering any more throughput or latency degradation than ICC under crash-faults. Finally, we want to show (iii) that Banyan is performing consistently as well as or better than ICC and other state-of-the-art protocols in a global network topology.
As the goal of our evaluation is to measure the precise impact of our changes, we define latency as the average proposal finalization time, measured at the respective proposer using their system clocks. In turn, throughput is measured as the average number of committed bytes per second at any (non-faulty) replica. We measure the protocols’ behavior under varying load by controlling the size of the payload, where the payload is a bit vector randomly generated by the leader.
@@
As required by the protocol, we set the proposal ($\Delta_{prop}$) and notarization ($\Delta_{notary}$) delays for Banyan and ICC to be larger than the message delay experienced without network disruptions, such that our experimental results correspond to regular network conditions and only one block is proposed per round if there are no faults. We proceed in the same way to set appropriate timeouts for HotStuff and Streamlet. Each experiment is run for 120 seconds, which is shown to be sufficient, as the measurements show remarkable regularity (e.g., see Figure 6(c)).
(a) Throughput vs. proposal latency for n=19 replicas spread across 4 global datacenters.
(b) Throughput vs. proposal latency for n=4 replicas spread across 4 global datacenters.
(c) Variance of Banyan and ICC proposal latencies compared with 1 MB payload and n=4.
(d) Effect of crash-faults on throughput and block intervals for n=19 replicas, spread across 4 US datacenters.
(e) Throughput vs. proposal latency for n=19 replicas spread across a global network.
Figure 6. Evaluation results.
9.3. Performance Evaluation
We distribute 19 replicas across 4 datacenters, as shown by the red triangles in Figure 5. Each datacenter hosts 5 replicas, except for one that hosts only 4. Figure 6(a) shows the throughput with varying block sizes. Primarily, we note that Banyan with $p=1$ performs consistently better or as well as ICC. For blocks of size 400KB, ICC averages a proposal finalization time of 239ms, while Banyan $p=1$ averages a finalization time of 216ms. This improvement of about $10\%$ is expected to be less than the theoretical maximum of $33\%$, because Banyan must hear from all the data centers in the fast path. In the slow path, replicas can collect two consecutive quorums without needing to communicate with the furthest datacenter. This observation, namely that more communication rounds with smaller quorum sizes are sometimes faster in practice than fewer rounds with larger quorum sizes, has been made before in similar contexts  (Junqueira et al., 2007; Sousa and Bessani, 2015). Moreover, by setting $p=f=4$, we observe that Banyan now has an average proposal finalization time of 179ms at the same block size, an improvement of 25.1%, that is much closer to the theoretical maximum of 33%. We suggest that this is due to the situation where the 4 co-located replicas are the furthest, in which case the fast path is employed, and thus lowers the average latency.
Next, we measure the potential of the Banyan fast path with a low number of replicas, i.e., we use only a single replica at each datacenter ($n=4$). Crucially, in this case the fast path fires with the same conditions as regular notarization, i.e., after receiving just 3 replies. Figure 6(b) shows the throughput with block size increments of 500KB. At block sizes of 1MB, ICC averages a proposal finalization time of 224ms, which Banyan manages to improve by 29.9% to 157ms. We emphasize that this large performance increase does not come at the cost of higher variance in latency, as is shown in Figure 6(c).
9.4. Effect of Crash-faults
In this experiment, we show the effect that crashes have on Banyan and Internet Computer Consensus. We distribute 19 replicas across four US AWS datacenters, as shown by the yellow crosses in Figure 5. Results are presented in Figure 6(d). As has been noted in literature, rotating leader protocols are sensitive to crashes (or malicious behavior), as at least one full timeout duration must expire before progress can be made in the case of a faulty leader. The timeout parameter will crucially impact the results of such scenarios. In the presented experiment, the timeout was set to 3 seconds. As seen in Figure 6(d), there are no penalties in trying to take the fast path. When they are failures, the performance of Banyan is exactly the one of ICC.
@@ 9.5. Global Network
In this experiment, we simulate a worldwide deployment. We include almost all AWS datacenter locations available to us and distribute 19 replicas across the globe, as shown by the black circles in Figure 5. In this setting and with 1MB payloads, the average proposal finalization time measured for the ICC implementation was 384ms. By running Banyan with $f=6,p=1$, the average proposal finalization time is reduced by 5.8% to 362ms for “free”. Further, for Banyan with $f=4,p=4$, the number drops further by 16% to 324ms (Figure 6(e)).
@@ 10. Conclusion
The proposed Banyan protocol improves upon current BFT protocols by providing a simultaneous dual mode, without incurring extra cost, thus closing the gap between state-of-the-art SMR protocols and classical fast consensus literature.
We show that it is possible to achieve optimally fast proposal finalization time for a rotating leader SMR protocol. We experimentally demonstrate the effectiveness of the fast path by showing consistent latency improvements compared to ICC, HotStuff and Streamlet. Ultimately, the effectiveness of the fast path depends largely on the network topology, as well as the parameter $p\in[0,f]$. Still, as we can set $p=1$ at essentially no cost, we hypothesize that Banyan can have large positive impact in practice, especially in networks with fewer replicas. In the context of permissioned blockchains, and future decentralized Layer-2 sequencers, this setting is undoubtedly compelling, since it is possible to detect adversarial behavior and remove misbehaving replicas (Buterin and Griffith, 2017; Camenisch et al., 2022).
11. Acknowledgements
We thank the anonymous Middleware reviewers for their extensive and valuable reviews, that among other things lead to the improvement of the resilience of Banyan. We further extend our gratitude to Prof. Christian Cachin for his insightful feedback.